OpenStudy (anonymous):

Integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ \sin^2(2x) }dx\]

OpenStudy (anonymous):

What is the integral of csc ^2 x ?

OpenStudy (anonymous):

@shailkumar it's not a usually one you don't need to know all the tables from universe

OpenStudy (anonymous):

if you are writing 1 as sin^2(x)+cos^2(x)=> \[x+\int\limits_{}^{}cotan^2(2x)\]

OpenStudy (anonymous):

Ok. But integral of csc^2 x, you should know. It is -cot x. Your integral is \[\int\limits \csc^2 2x dx\] = -cot 2x /2

OpenStudy (anonymous):

Of course an additive constant.

OpenStudy (anonymous):

it's not in my tables but I've kind found a way to solve it watch it \[\sin(2x)=2sinxcosx=>\int\limits_{}^{}(2\sin(x)\cos(x))^{-2} dx\]

OpenStudy (anonymous):

now if we substitute sin(x)=t=>cosxdx=dt

OpenStudy (anonymous):

so far good, then ?

OpenStudy (anonymous):

You will be stuck to handle the swinging -2.

OpenStudy (anonymous):

Ok. what is the derivative of cot x ?

OpenStudy (jmark):

Integ (1/sin^2 (2x) dx = -1/2 cot (2x) +c Get ICSE study material - http://www.pinterest.com/tramesh/icse/